# Mitigating Multimodal Hallucinations via Gradient-based Self-Reflection

### v3

@@ 5.4 Gradient Computation and Norm Analysis
Our method obtains gradients directly through PyTorch’s ‘torch.autograd.grad’ on input tokens, eliminating the need for manual derivations and enabling straightforward reproducibility. For comparison, we evaluate Integrated Gradients (IG) [enguehard2023sequential, lundstrom2022rigorous, kapishnikov2021guided] using the SIG [enguehard2023sequential] implementation; results for this ablation on the POPE MSCOCO adversarial setting are shown in Tab. 6. ‘IG’ denotes the SIG-based results, while “Ours” refers to our direct gradient method. Both achieve comparable accuracy and F1‐score, but the direct-gradient variant is substantially more efficient.
@@ 5 Results on MMBench
We further evaluate our method on MMBench [liu2025mmbench]. The results in Tab. 3 indicate that our method improves the overall performance and achieves consistent improvements across MLLMs on Coarse Perception (CP). This outcome aligns with the intended effect of our method, as its focus on increasing visual influence is directly linked to improving coarse perception capabilities. For other metrics, our method yields minor improvements due to the possible reason that certain abilities, such as Logical Reasoning (LR), rely more on the language component of MLLMs and cannot be enhanced solely by increasing visual influence.
@@ 6 Results on MM-Vet
The evaluation on MM-Vet [yu2023mm] in Tab. 4 shows that our method achieves consistent overall (Total) improvement, along with enhancements in recognition (Rec) and Optical Character Recognition (OCR), indicating its effectiveness in improving visual recognition. However, its performance varies across other metrics, including knowledge (Know), generalization (Gen), spatial awareness (Spat), and math (Math), suggesting that our method, which focuses on token influence balancing, may not effectively enhance the generalization ability of MLLMs.
@@ 7 Results on ScienceQA and Vizwiz
We evaluate our method on two complementary multimodal benchmarks. ScienceQA [lu2022learn] integrates images, textual context, and curriculum knowledge, requiring models to perform structured multimodal reasoning. VizWiz [gurari2018vizwiz], in contrast, consists of visual questions collected from blind users and features real-world challenges such as low-quality images, conversational queries, and unanswerable cases. These datasets jointly assess both reasoning under structured multimodal contexts and robustness in unconstrained real-world settings. As shown in Table 5, our approach consistently improves over the LLaVA-1.5 baseline. These gains demonstrate the effectiveness of our hallucination mitigation strategy in enhancing visual grounding across both knowledge-driven and real-world VQA tasks.
@@ 8 Results on MME
Our evaluation on MME [fu2024mmecomprehensiveevaluationbenchmark] dataset is presented in Tab. 6. Our method achieves better overall (Total) results and equal or improved performance in existence and counting, demonstrating its effectiveness in object recognition. However, it does not improve position accuracy and exhibits varying behavior on color. This diversity may stem from the inherent capabilities of MLLMs, which cannot be solely enhanced through token influence balancing.
@@ 9 Other Results of POPE
Table 7: More Results on POPE [li2023evaluating].
@@
MSCOCO Random base 86.5 84.8 87.1 85.7 86.0 84.4
ours 86.8 85.1 87.9 86.8 87.9 87.1
Popular base 85.5 83.8 84.2 83.6 84.6 83.2
ours 85.6 84.0 85.0 84.3 86.4 85.7
@@ 10 Different Sampling Strategies
Tab. 8 presents an ablation study on sampling strategies (non-greedy vs. greedy). We follow the non-greedy sampling setting of VCD [leng2024mitigating], where both top-p and temperature are set to 1. As shown, our method consistently improves performance across both sampling strategies.
@@ 13 Grouping Multiple Influential Tokens with Respect to the Anchor Object
We also explored top-$2$ variants as a straightforward extension, but they produced more diffuse attributions and did not yield consistent improvements over the top-$1$ design.
@@ 14 Hyper parameter Study
Maximum $\alpha_{m}$. To determine the optimal maximum amplification factor ($\alpha_{m}$) and understand its impact on the model performance, we conducted a search over values from 1 to 6 using the LLaVA-v1.5 model on the POPE dataset for discriminative tasks. For open-ended generation tasks on a subset of the MSCOCO dataset following [deng2024seeing], we observed garbled text when $\alpha_{m}$ was set to 5; therefore, we limited our search to values from 1 to 4. As shown in Tab. 12 and Tab. 13, discriminative task on POPE is less sensitive to the value of $\alpha_{m}$. The performance on the generative task gets improved while $\alpha_{m}$ increases but later drops. Therefore, in our experiments, the optimal maximum amplification factor ($\alpha_{m}$) is set to 5 and 3 for discriminative and generation tasks, respectively.
Table 12: $\alpha_{m}$ Study For Discriminative Task On POPE [li2023evaluating] in MSCOCO Adversarial Setting.
@@
Early Stopping Threshold. To set the early stopping threshold properly, we conducted a search on a subset of the MSCOCO dataset subset following [deng2024seeing]. Recognizing that the visual influence ratio varies across models, we first analyze the sample-wise visual influence to identify an appropriate range for this study. By searching over the corresponding range, we show results in Tab. 14. These results demonstrate that varying the ES threshold primarily mediates the trade-off between recall and hallucination rate. Our goal is to have balanced recall ($R$) and instance-level hallucination ($C_{I}$), leading us to select thresholds of $7\%$ for LLaVA-v1.5 and LLaVA-v1.6, $25\%$ for InstructBLIP, $2.5\%$ for mPLUG-Owl2 and $10\%$ for InternVL2. We additionally ran an experiment to measure the ES activation rate using LLaVA-v1.5 with ES threshold $7\%$. As shown in Tab. 15, ES fires on only $8.7\%$ of the test samples, and when it does, the generated responses are on average just 0.7 tokens shorter. This indicates that ES rarely, and only minimally, truncates outputs.
@@ 19 Qualitative Example on Occluded Images
20 Additional Implementation and Experimental Details
Noun tokens are identified using the spaCy library via its en_core_web_sm model. For experiments on the Amber dataset [wang2023llm], we adopt the original data splits and evaluation metrics. In the MSCOCO [lin2014microsoft] subset, we follow the data partitioning and evaluation protocol of Deng et al. [deng2024seeing], with splits available in their official repository. For the LLaVA-QA90 [liu2024visual], MME [fu2024mmecomprehensiveevaluationbenchmark] and POPE [li2023evaluating] datasets, our setup replicates that of Leng et al. [leng2024mitigating] and use their provided scoring scripts for LLaVA-QA90. Experiments on MMBench [liu2025mmbench], MM-Vet [yu2023mm] follow the VLMEvalKit_InternVL2_5 repository. All comparison methods are executed using their official code; we only modify them to enforce greedy sampling and a uniform maximum generation length to align with our experimental settings.
21 Computational Cost
In this section, we analyze the computational cost of our method. The visual encoder is executed only once, and the second pass operates on a small set of tokens. On the POPE dataset, this results in a $101.44\%$ increase in average TFLOPs, which is consistent with decoding-based approaches that rely on two forward passes. For comparison, VCD incurs an overhead of $100.10\%$.
Method TFLOPs Relative Increase
LLaVA-v1.5 $9.68$
VCD $19.37$ $100.10\%$
Ours $19.49$ $101.44\%$
22 Broader Impacts

### v4

@@ 5.4 Gradient Computation and Norm Analysis
Our method obtains gradients directly through PyTorch’s ‘torch.autograd.grad’ on input tokens, eliminating the need for manual derivations and enabling straightforward reproducibility. For comparison, we evaluate Integrated Gradients (IG) [11, 36, 23] using the SIG [11] implementation; results for this ablation on the POPE MSCOCO adversarial setting are shown in Tab. 6. ‘IG’ denotes the SIG-based results, while “Ours” refers to our direct gradient method. Both achieve comparable accuracy and F1‐score, but the direct-gradient variant is substantially more efficient.
@@ 2 Interpreting Contrastive Decoding through KL Divergence
3 Derivation of $\alpha$ Computation
We choose $\alpha_{m}$ such that the influence of $\mathbf{t}^{u}$ matches the dominant text level. For clarity, we present the derivation for the case where the prompt is dominant; the case where the previous output tokens are dominant is analogous and leads to the same derivative up to a straightforward substitution. Given $\hat{\texttt{I}}^{v}_{m}=(1+\alpha_{m})\texttt{I}^{v}_{m}-\alpha_{m}\tilde{\texttt{I}}^{o}_{m}$ and $\hat{\texttt{I}}^{p}_{m}=(1+\alpha_{m})\texttt{I}^{p}_{m}-\alpha_{m}\tilde{\texttt{I}}^{p}_{m}$, we enforce $\hat{\texttt{I}}^{v}_{m}=\hat{\texttt{I}}^{p}_{m}$, and solve for $\alpha_{m}$:
$(1+\alpha_{m})\texttt{I}^{v}_{m}-\alpha_{m}\tilde{\texttt{I}}^{o}_{m}$ $=(1+\alpha_{m})\texttt{I}^{p}_{m}-\alpha_{m}\tilde{\texttt{I}}^{p}_{m}$ (18)
$\alpha_{m}(\texttt{I}^{v}_{m}-\tilde{\texttt{I}_{m}}^{o}+\tilde{\texttt{I}}^{p}_{m}-\texttt{I}^{p}_{m})$ $=\texttt{I}^{p}_{m}-\texttt{I}^{v}_{m}$
$\alpha_{m}$ $=\frac{\texttt{I}^{p}_{m}-\texttt{I}^{v}_{m}}{\texttt{I}^{v}_{m}-\tilde{\texttt{I}_{m}}^{o}+\tilde{\texttt{I}}^{p}_{m}-\texttt{I}^{p}_{m}}.$
@@ 5 MLLMs Architectures
Qwen2-VL-7B QwenViT Qwen2-7B
Qwen2-VL-7B QwenViT Qwen2.5-7B
@@ 6 Results on MMBench
We further evaluate our method on MMBench [34]. The results in Tab. 3 indicate that our method improves the overall performance and achieves consistent improvements across MLLMs on Coarse Perception (CP). This outcome aligns with the intended effect of our method, as its focus on increasing visual influence is directly linked to improving coarse perception capabilities. For other metrics, our method yields minor improvements due to the possible reason that certain abilities, such as Logical Reasoning (LR), rely more on the language component of MLLMs and cannot be enhanced solely by increasing visual influence.
@@ 7 Results on MM-Vet
The evaluation on MM-Vet [53] in Tab. 4 shows that our method achieves consistent overall (Total) improvement, along with enhancements in recognition (Rec) and Optical Character Recognition (OCR), indicating its effectiveness in improving visual recognition. However, its performance varies across other metrics, including knowledge (Know), generalization (Gen), spatial awareness (Spat), and math (Math), suggesting that our method, which focuses on token influence balancing, may not effectively enhance the generalization ability of MLLMs.
@@ 8 Results on ScienceQA and Vizwiz
We evaluate our method on two complementary multimodal benchmarks. ScienceQA [35] integrates images, textual context, and curriculum knowledge, requiring models to perform structured multimodal reasoning. VizWiz [16], in contrast, consists of visual questions collected from blind users and features real-world challenges such as low-quality images, conversational queries, and unanswerable cases. These datasets jointly assess both reasoning under structured multimodal contexts and robustness in unconstrained real-world settings. As shown in Table 5, our approach consistently improves over the LLaVA-1.5 baseline. These gains demonstrate the effectiveness of our hallucination mitigation strategy in enhancing visual grounding across both knowledge-driven and real-world VQA tasks.
@@ 9 Results on MME
Our evaluation on MME [13] dataset is presented in Tab. 6. Our method achieves better overall (Total) results and equal or improved performance in existence and counting, demonstrating its effectiveness in object recognition. However, it does not improve position accuracy and exhibits varying behavior on color. This diversity may stem from the inherent capabilities of MLLMs, which cannot be solely enhanced through token influence balancing.
@@ 10 Other Results of POPE
Table 7: More Results on POPE [29].
@@
MSCOCO Random base 87.1 85.4 87.1 85.7 86.0 84.4
ours 87.4 86.0 87.9 86.8 87.9 87.1
Popular base 85.9 84.4 84.2 83.6 84.6 83.2
ours 86.2 84.8 85.0 84.3 86.4 85.7
@@ 11 Different Sampling Strategies
Tab. 8 presents an ablation study on sampling strategies (non-greedy vs. greedy). We follow the non-greedy sampling setting of VCD [26], where both top-p and temperature are set to 1. As shown, our method consistently improves performance across both sampling strategies.
@@ 14 Grouping Multiple Influential Tokens with Respect to the Anchor Object
We also explored top-$2$ variants as a straightforward extension, but they did not yield consistent improvements over the top-$1$ design. One possible reason is that, for some predicted tokens, supervision is effectively dominated by a single visual token, so forcing a top-$2$ aggregation can dilute this primary contribution rather than help. As a result, the top-$2$ scheme may only benefit a subset of objects that are genuinely supported by multiple visual tokens, leading to limited overall gains.
@@ 15 Hyper parameter Study
Maximum $\alpha_{m}$ serves primarily as a precautionary upper bound. Because Eq. 13 in the main text naturally bounds $\alpha$, this maximum threshold is rarely triggered, as evidenced by the low clipping rates reported in Tab. 12.
Table 12: Clipping Rate of $\alpha_{m}$ at Maximum Value
Task Discriminative @ POPE Adversarial Setting Generation @ MSCOCO Subset
GACD (LLaVA-v1.5) 0.04% 0.07%
Nevertheless, to determine the optimal value for $\alpha_{m}$ and assess its impact on model performance, we conducted a hyperparameter search using LLaVA-v1.5. For discriminative tasks, we evaluated $\alpha_{m}\in\{1,\dots,6\}$ on the POPE dataset. For open-ended generation tasks on a subset of MSCOCO [10], we observed garbled text outputs when $\alpha_{m}\geq 5$; thus, we restricted our search space to $\alpha_{m}\in\{1,\dots,4\}$. As shown in Tab. 13 and Tab. 14, performance on the POPE discriminative task is relatively insensitive to variations in $\alpha_{m}$. Conversely, performance on the generative task initially improves as $\alpha_{m}$ increases but degrades at higher values. Based on these findings, we set the optimal maximum amplification factor $\alpha_{m}$ to 5 for discriminative tasks and 3 for generative tasks across all our experiments.
Table 13: $\alpha_{m}$ Study For Discriminative Task On POPE [29] in MSCOCO Adversarial Setting.
@@
Early Stopping Threshold. To set the early stopping threshold properly, we conducted a search on a subset of the MSCOCO dataset subset following [10]. Recognizing that the visual influence ratio varies across models, we first analyze the sample-wise visual influence to identify an appropriate range for this study. By searching over the corresponding range, we show results in Tab. 15. These results demonstrate that varying the ES threshold primarily mediates the trade-off between recall and hallucination rate. Our goal is to have balanced recall ($R$) and instance-level hallucination ($C_{I}$), leading us to select thresholds of $7\%$ for LLaVA-v1.5 and LLaVA-v1.6, $25\%$ for InstructBLIP, $2.5\%$ for mPLUG-Owl2 and $10\%$ for InternVL2. We additionally ran an experiment to measure the ES activation rate using LLaVA-v1.5 with ES threshold $7\%$. As shown in Tab. 16, ES fires on only $8.7\%$ of the test samples, and when it does, the generated responses are on average just 0.7 tokens shorter. This indicates that ES rarely, and only minimally, truncates outputs.
@@ 16 Confidence and Visual Influence
17 Additional Implementation and Experimental Details
We identify noun tokens using the spaCy library due to its lightweight, CPU-only operation; this component is interchangeable with any alternative noun detector. At each step, if the current token is classified as a noun and at least one noun has appeared in previous outputs, we trigger noun-only grouping. In this mode, we inject noun-related visual tokens into the negative guidance; otherwise, negative guidance is computed using only text tokens. For experiments on the Amber dataset [48], we adopt the original data splits and evaluation metrics. In the MSCOCO [30] subset, we follow the data partitioning and evaluation protocol of Deng et al. [10], with splits available in their official repository. For the LLaVA-QA90 [32], MME [13] and POPE [29] datasets, our setup replicates that of Leng et al. [26] and use their provided scoring scripts for LLaVA-QA90. Experiments on MMBench [34], MM-Vet [53] follow the VLMEvalKit_InternVL2_5 repository. All comparison methods are executed using their official code; we only modify them to enforce greedy sampling and a uniform maximum generation length to align with our experimental settings.
18 Computational Cost
Method TFLOPs Runtime (ms) Relative Increase
LLaVA-v1.5 $9.68$ 191
VCD $19.37$ 383 $100.10\%$
Ours $19.49$ 385 $101.44\%$
In this section, we analyze the computational cost of our method. The visual encoder is executed only once, and the second pass operates on a small set of tokens. On the POPE dataset, this results in a $101.44\%$ increase in average TFLOPs/runtime, which is consistent with decoding-based approaches that rely on two forward passes. For comparison, VCD incurs an overhead of $100.10\%$.
19 Broader Impacts